Find the number of neutrons released by the following fission reaction.

1
0

n +

235
92

U ?  

117
50

Sn +

101
42

Mo +  neutron(s)
Answer #1

on the reactants side ::

total number of neutrons = 1 + (235 - 92) = 144

on the product side, number of neutrons = n + (117 - 50) + (101 - 42)

ideally

neutrons on reactants side = neutrons on product side

144 = n + (117 - 50) + (101 - 42)

n = 18